OpenStudy (anonymous):

Solve 3x - 3 = x + 7

OpenStudy (anonymous):

3x - 3 = x + 7 2x - 3 = 7 (subtract x from both sides) 2x = 10 (add 3 to both sides) x = 5 (divide by 2)

OpenStudy (anonymous):

thank you